OpenStudy (anonymous):

Solve for the variable in this equation. 5x/3+2/7=x/9-4/7

OpenStudy (inkyvoyd):

Try multipling both sides by nine, and tell me if you get it.

Parth (parthkohli):

35x + 6 7x - 36 -------- = --------- 21 63 Cross multiply and solve it further :)

OpenStudy (anonymous):

would it be -55 over 96?

jhonyy9 (jhonyy9):

first group terms by terms 5x/3 -x/9 =-2/7 -4/7 (5x*3 -x)/9 = -6/7 (15x-x)/9 = -6/7 14x/9 = -6/7 7*14x/63=-6*9/63 7*14x= -6*9 98x=-54 49x=-27 x= -27/49

OpenStudy (anonymous):

Thanks so much:)

jhonyy9 (jhonyy9):

yw good luck bye